Hi and thank you for answering my question:
If 1/a = 1/b + 1/c, find the value of a where b= -4 and c= 6.5.
Thanks! :)

Answer by stanbon(75887)

If 1/a = 1/b + 1/c, find the value of a where b= -4 and c= 6.5.
-------
1st:: Solve for "a":
1/a = (c+b)/bc
-----------------
a = (bc)/(b+c)
-------
Evaluate::
a = (-4*6.5)/(-4+6.5)
----
a = (-26)/2.5
-----
a = -10.4
--------------
Cheers,
Stan H.
-------------